# Token-based Audio Inpainting via Discrete Diffusion

### v3

@@ 4.3 Span Masking
Whereas prior discrete diffusion frameworks (lou2310discrete, ) apply masking independently in the forward process, we propose a structured corruption strategy that enables the model to better handle inpainting. We employ a span-based masking strategy for the absorbing diffusion process.
Given the sequence of tokens $X=(x_{1},\dots,x_{L})$ of length $L$, we construct a subset of perturbed tokens by iteratively sampling contiguous spans until a predefined budget is met. The budget corresponds to the expected number of tokens transitioned away from their original state, defined as $B(t)=\big(1-e^{-\sigma(t)}\big)\cdot L$, where $\sigma(t)$ denotes the noise schedule at time $t$, and $1-e^{-\sigma(t)}$ represents the probability of transitioning away from the current state. This formulation ensures that the proportion of perturbed tokens aligns with the transition probability induced by the forward diffusion process.
@@ 5.1 Objective Metrics
We use FAD (kilgour2018fr, ), which measures the distributional distance between real and generated audio features. It reflects perceptual similarity in attributes such as timbre and texture, serving as an indicator of realism. Our evaluation uses a VGG-based feature extractor.
@@
Perceptual quality is further assessed with ODG from the PEMO-Q model (huber2006pemo, ), which simulates auditory processing and maps a Perceptual Similarity Measure (PSM) to scores from 0 (imperceptible) to –4 (very annoying). We also assessed with ODG using PEA-Q. Both were included to match the evaluation methods of related works. As a reference-based metric aligned with listening tests, ODG complements waveform-based measures.
@@
Following (moliner2023diffusion, ), we compute LSD to capture spectral distortions:
@@ 5.2 Experimental setup
We evaluated our model on two benchmark datasets. First, we used the MusicNet dataset (dataset, ), which provides a predefined split of training and test sets. The training set comprises 330 freely licensed classical music recordings with aligned labels for instruments, notes, and other musical attributes. This dataset serves as a compact benchmark for evaluating models on polyphonic music transcription. We report results using the predefined split.
In addition, we evaluated our model on the MAESTRO dataset (MAESTRO, ), which also provides predefined training and test splits. MAESTRO contains over 200 hours of virtuosic piano performances with precise note-level alignment between audio and MIDI. Due to its large size, we trained only on a subset of the official training set (see Supplementary Material for details).
We conducted experiments on each dataset independently; there was no mixing or combined training across datasets.
@@
During training, raw audio signals are first tokenized and truncated to a fixed length of 1024 tokens (approximately 13.6 seconds). A random timestep is selected, and noise (in the form of token masking) is applied corresponding to the span-based masking corruption. The DDM is then trained to predict the concrete scores from this corrupted input. Through reverse diffusion steps, the model gradually denoises the token sequence, effectively learning to recover clean representations from noisy ones in token space, illustrated in Figure 1.
Training employed the AdamW optimizer (learning rate of $10^{-6}$), batch size 128, and 1024 token samples. On MusicNet, the base model with DWDSE loss was trained for 400k steps (two days) on a single NVIDIA A6000 GPU, while other methods were trained for 100k steps. On MAESTRO, training ran for 150k steps (24h) under the same setup. Full hyperparameters are reported in the supplementary, following ablation results (see Section 7).
@@
We evaluate the effectiveness of our proposed audio inpainting method, we followed the experimental protocol of (moliner2023diffusion, ) and selected 60 previously unseen music segments from the MusicNet test set (dataset, ), each with a duration of 4.17 seconds. For each segment, we introduced four synthetic gaps at fixed, evenly spaced locations. The gap durations varied across experiments, ranging from 150 ms to 300 ms. Since AIDD is stochastic, we generated 10 samples for each segment and reported the averaged score as the final result.
For the MAESTRO dataset (MAESTRO, ), we followed (aironi2023time, )’s experimental set-up, taking 6 seconds segments from the predefined test set and introducing single gaps in the middle.
@@
At inference, the masked waveform is tokenized with the pretrained WavTokenizer encoder. The trained DDM then performs reverse diffusion to predict missing tokens conditioned on context. The token sequence is decoded back to waveform, and only the inpainted segment replaces the original gap, avoiding unnecessary re-encoding of intact regions. A short 6 ms crossfade is applied at boundaries to ensure smooth transitions. An overview is shown in Figure 1.
@@ 6 Results
For the MusicNet dataset, we evaluate our method against three baselines: LPC, an extrapolation method using linear predictive coding (kauppinen2002audio, ); A-SPAIN-L, a sparsity-based inpainting approach with dictionary learning (taubock2020dictionary, ); and CQT-Diff+, a diffusion model with Constant-Q Transform, state-of-the-art for gaps up to 300 ms (moliner2023diffusion, ).
For the MAESTRO dataset (MAESTRO, ) we evaluated against three baselines as well: GACELA (marafioti2020gacela, ): a cGAN that performs long-gap audio inpainting using multiple discriminators at different time scales; bin2bin (aironi2023score, ): a spectrogram inpainting model based on a pix2pix-style GAN with STFT losses; bin2bin-MIDI (aironi2023time, ): an extension of bin2bin that adds a piano-roll (MIDI) loss to enforce pitch consistency.
Table 1 reports the average performance across varying gap durations on the MusicNet (dataset, ) dataset, while Table 2 presents the corresponding results on the MAESTRO (MAESTRO, ) dataset.
@@
Table 1: Performance comparison of methods across different gap lengths using FAD, LSD, and ODG metrics, on MusicNet dataset (dataset, ). Lower is better for FAD/LSD ($\downarrow$), higher is better for ODG ($\uparrow$). In bold the best score. The AIDD presented is the highest reported scores, as described in section 5.2.
@@
AIDD -2.370 $\pm$ 0.705 -2.959 $\pm$ 0.604
Table 2: ODG (PEA-Q) score values for the three compared methods to AIDD, on MAESTRO (MAESTRO, ) dataset. In bold the best score. Higher is better ($\uparrow$). The AIDD’s result reported is the combined method.
On the MusicNet dataset, our method (AIDD) consistently outperforms prior approaches across most gap durations. In particular, AIDD achieves the lowest FAD values for gaps of 200-300 ms, while also obtaining the best LSD and generally stronger ODG scores. Compared to the strong diffusion-based baseline CQT-Diff+, our approach yields significantly lower distortion (FAD/LSD) for medium and longer gaps, with improvements of nearly 25% in FAD at 300 ms (3.549 vs. 4.652). This suggests that AIDD generates more perceptually consistent reconstructions, especially as the gap size increases. At short gaps (150 ms), CQT-Diff+ achieves a lower FAD, though the margin is small; nevertheless, AIDD provides better ODG and LSD in this setting. Finally, it is worth noting that our base model is half the size of CQT-Diff+ and required only one-quarter of the training time, underscoring its efficiency.
On the MAESTRO dataset, results follow a similar trend (Table 2). Our method (AIDD) achieves the best ODG score for the 375 ms gap, outperforming all three baselines by a clear margin (2.370-2.370 vs. 2.800-2.800 to 3.232-3.232). This indicates that AIDD can better preserve perceptual quality even in the presence of interruptions. At 750 ms, our model also achieves the top performance (2.959-2.959), surpassing all three baselines, which was specifically designed for long-gap inpainting.
It is important to highlight that, unlike prior works which trained on the full MAESTRO dataset, we only used a reduced training subset, yet still achieved state-of-the-art performance for both shorter (375 ms) and longer (750 ms) gaps. This suggests that AIDD generalizes well to long-context musical structures while requiring fewer resources, both in terms of dataset size and computational budget.
@@ 7 Ablation Study
Table 3: Performance comparison of different AIDD methods across different gap lengths using FAD, LSD, and ODG metrics, on MusicNet dataset. In bold the best scores. Lower is better for FAD/LSD ($\downarrow$), higher is better for ODG ($\uparrow$).
@@ 8 Conclusion
In this work, we proposed a novel approach for audio inpainting by combining a DDM with a WavTokenizer-based encoder-decoder architecture (ji2024wavtokenizer, ). Rather than applying diffusion directly to raw waveforms, which are noisy and difficult to model, we operate in a discrete token space, enabling the DDM to effectively learn the underlying distribution of musical structure. This allows for more stable training and more coherent reconstruction during inference. Our method not only advances the task of audio inpainting, but also lays the foundation for additional generative capabilities, including unconditional music generation and music continuation based on a given context. Moreover, we introduce two novel training methods for the discrete diffusion process: span-based masking and derivative-based loss that better contribute to the model’s learning of both fine-grained and structured corruptions.
Experimentally, we demonstrated that our method achieves competitive or superior results compared to existing state-of-the-art approaches, particularly in challenging settings with longer gaps. While many previous methods focus on short gaps of up to 100 ms, our model maintains strong performance even on 750 ms gaps, highlighting its potential applicability to real-world scenarios where such gaps can significantly affect the listening experience. Overall, our method introduces a robust and extensible paradigm for music-based generative modeling in the token domain, with promising applications in both restoration and creative audio generation tasks, as well as lays the foundation for continuous work in the field of DDM training.
While our study focused on music inpainting, the proposed architecture is not limited to this domain. Future research will be extending the framework to speech audio inpainting, where missing segments can disrupt the intelligibility of entire sentences. Addressing such scenarios would demonstrate the broader applicability of discrete diffusion for real-world communication, and position this line of work as a practical solution for diverse audio restoration tasks.
Future work may extend our training strategies to large language models by building on our base method. Although audio and language differ inherently, both can be approached through token-based discrete diffusion.
@@
Appendix A Code for AIDD
Our full code will be published in GitHub for open access once the paper is accepted. Currently, the code is added as a folder with the submission.
@@ Appendix B Reproducibility Statement
With the following hyperparameter as well as appropriate hardware presented in our supplementary material, anyone can reproducibly conduct our experiment using our code which, upon acceptance, will be publicly accessed.
Appendix C Reproducibility Materials
C.1 Hyperparameters
@@
Sampling rate 128
@@
Derivative $\Delta^{1}e$
@@ C.2 Training Environment
Appendix D MAESTRO dataset
Due to its large size, we randomly took a subset of the training and test, following the test-train split predefined. Upon request, we will provide the full list.

### v4

@@ 4.3 Span Masking
Whereas prior discrete diffusion frameworks (Lou et al., 2024) apply masking independently in the forward process, we propose a structured corruption strategy that enables the model to better handle inpainting. We employ a span-based masking strategy for the absorbing diffusion process.
Given the sequence of tokens $X=(x_{1},\dots,x_{L})$ of length $L$, we construct a subset of perturbed tokens by iteratively sampling contiguous spans until a predefined budget is met. This iterative procedure may produce multiple disjoint masked spans within a single timestep. The budget corresponds to the expected number of tokens transitioned away from their original state, defined as $B(t)=\big(1-e^{-\sigma(t)}\big)\cdot L$, where $\sigma(t)$ denotes the noise schedule at time $t$, and $1-e^{-\sigma(t)}$ represents the probability of transitioning away from the current state. This formulation ensures that the proportion of perturbed tokens aligns with the transition probability induced by the forward diffusion process.
@@ 5.1 Objective Metrics
We use FAD (Kilgour et al., 2018), which measures the distributional distance between real and generated audio features. It reflects perceptual similarity in attributes such as timbre and texture, serving as an indicator of realism. Our evaluation uses a VGG-based feature extractor.
@@
Perceptual quality is further assessed with ODG from the PEMO-Q model (Huber and Kollmeier, 2006), which simulates auditory processing and maps a Perceptual Similarity Measure (PSM) to scores from 0 (imperceptible) to –4 (very annoying). We also assessed with ODG using PEA-Q. Both were included to match the evaluation methods of related works. As a reference-based metric aligned with listening tests, ODG complements waveform-based measures.
@@
Following (Moliner and Välimäki, 2023), we compute LSD to capture spectral distortions:
@@
Subjective Listening test.
We conduct a mean opinion score (MOS) test to assess perceptual quality. Participants rate randomly ordered audio segments (1-5 scale), covering different gap lengths and anonymized across methods; original segments are included as a reference. Full evaluation details appear in the supplementary material.
@@ 5.2 Experimental setup
We evaluated our model on two benchmark datasets: MusicNet (Thickstun et al., 2016) and MAESTRO (Hawthorne et al., 2019). MusicNet contains 330 classical music recordings with aligned instrument and note annotations. MAESTRO includes over 200 hours of aligned piano performances. We used the predefined train/test splits and trained on each dataset independently.
@@
During training, raw audio signals are first tokenized and truncated to a fixed length of 300 tokens (approximately 4 seconds). A random timestep is selected, and noise (in the form of token masking) is applied corresponding to the span-based masking corruption. The DDM is then trained to predict the concrete scores from this corrupted input. Through reverse diffusion steps, the model gradually denoises the token sequence, effectively learning to recover clean representations from noisy ones in token space, illustrated in Figure 1.
Training employed the AdamW optimizer (learning rate of $10^{-6}$), batch size 128, and 300 token samples. On MusicNet, the base model with DWDSE loss was trained for 400k steps (two days) on a single NVIDIA A6000 GPU, while other methods were trained for 100k steps. On MAESTRO, training ran for 150k steps (24h) under the same setup. Full hyperparameters are reported in the supplementary, following ablation results (see Section 7).
@@
We evaluate the effectiveness of our proposed audio inpainting method, we followed the experimental protocol of (Moliner and Välimäki, 2023) and selected 60 previously unseen music segments from the MusicNet test set (Thickstun et al., 2016), each with a duration of 4.17 seconds. For each segment, we introduced four synthetic gaps at fixed, evenly spaced locations. The gap durations varied across experiments, ranging from 150 ms to 300 ms. Since AIDD is stochastic, we generated 10 samples for each segment and reported the averaged score as the final result.
For the MAESTRO dataset (Hawthorne et al., 2019), we followed (Aironi et al., 2023b)’s experimental set-up, taking 6 seconds segments from the predefined test set and introducing single gaps in the middle.
@@
For subjective evaluation, we introduce MOS subjective listening test, giving the same audio segments for different methods and asking people to rate the quality of audio without prior information on the method, gap or file name. We perform this on the MAESTRO dataset. (The full setup is in the Supplementary Materials).
@@
At inference, the masked waveform is tokenized with the pretrained WavTokenizer encoder. The trained DDM then performs reverse diffusion to predict missing tokens conditioned on context. The token sequence is decoded back to waveform, and only the inpainted segment replaces the original gap, avoiding unnecessary re-encoding of intact regions. A short 10 ms crossfade is applied at boundaries to ensure smooth transitions. An overview is shown in Figure 1.
@@ 6 Results
For the MusicNet dataset, we evaluate our method against three baselines: LPC, an extrapolation method using linear predictive coding (Kauppinen and Roth, 2002); A-SPAIN-L, a sparsity-based inpainting approach with dictionary learning (Tauböck et al., 2020); and CQT-Diff+, a diffusion model with Constant-Q Transform, state-of-the-art for gaps up to 300 ms (Moliner and Välimäki, 2023).
For the MAESTRO dataset (Hawthorne et al., 2019) we evaluated against three baselines as well: GACELA (Marafioti et al., 2020): a cGAN that performs long-gap audio inpainting using multiple discriminators at different time scales; bin2bin (Aironi et al., 2023a): a spectrogram inpainting model based on a pix2pix-style GAN with STFT losses; bin2bin-MIDI (Aironi et al., 2023a): an extension of bin2bin that adds a piano-roll (MIDI) loss to enforce pitch consistency.
Table 1 reports the average performance across varying gap durations on the MusicNet (Thickstun et al., 2016) dataset, while Table 2 presents the corresponding results on the MAESTRO dataset.
@@
Table 1: Performance comparison of methods across different gap lengths using FAD, LSD, and ODG metrics, on MusicNet dataset (Thickstun et al., 2016). Lower is better for FAD/LSD ($\downarrow$), higher is better for ODG ($\uparrow$). In bold the best score. The AIDD presented is the highest reported scores, as described in section 5.2.
@@
AIDD -2.303 $\pm$ 0.692 -2.596 $\pm$ 1.300
Table 2: ODG (PEA-Q) score values for the three compared methods to AIDD, on MAESTRO (Hawthorne et al., 2019) dataset. In bold the best score. Higher is better ($\uparrow$). The AIDD’s result reported is the combined method.
On the MusicNet dataset, AIDD outperforms prior approaches across most gap durations. It achieves the lowest FAD for 200–300 ms gaps and consistently improves LSD and ODG compared to baselines. Against the strong diffusion-based CQT-Diff+, AIDD yields substantially lower distortion for medium and long gaps, with a $\sim 25\%$ FAD reduction at 300 ms (3.549 vs. 4.652). This indicates more perceptually coherent reconstructions as gap size increases. For short gaps (150 ms), CQT-Diff+ attains slightly better FAD, but AIDD still delivers superior ODG and LSD.
Finally, it is worth noting that our base model is almost half the size of CQT-Diff+ and required only one-quarter of the training time, underscoring its efficiency (see Section 8 for latency analysis).
On the MAESTRO dataset, results follow a similar trend (Table 2). Our method (AIDD) achieves the best ODG score for the 375 ms gap, outperforming all three baselines by a clear margin. This indicates that AIDD can better preserve perceptual quality even in the presence of interruptions.
@@
Beyond excelling in objective metrics, AIDD continues to outperform baselines in subjective listening assessments, as shown in Table 3.
Another major advantage to AIDD is that out of all the methods we compared to, our method is the only one that does not require fixed size of input audio.
Method MOS ($\uparrow$)
Original 4.12 $\pm$ 0.96
GACELA 3.51 $\pm$ 1.33
CQT-Diff+ 3.51 $\pm$ 1.34
AIDD (WavTokenizer 24kHz) 3.64 $\pm$ 1.26
Table 3: MOS values for the compared methods to AIDD, on MAESTRO (Hawthorne et al., 2019) dataset. In bold the best score. Higher is better ($\uparrow$). The AIDD’s result reported is the combined method.
@@ 7 Ablation Study
We also compared between two different tokenizers: UniCodec (Jiang et al., 2025) and WavTokenizer (Ji et al., 2024) on MAESTRO dataset, results shown in Table 5. We also present information loss analysis study between the two in the Supplementary Material. We speculate that although Unicodec pressumed to have better audio quality, the larger codec ($\sim$16k) vs WavTokenizer’s ($\sim$4k) may be too large for our model.
@@
Table 4: Performance comparison of different AIDD methods with WavTokenizer, across different gap lengths using FAD, LSD, and ODG metrics, on MusicNet dataset. In bold the best scores. Lower is better for FAD/LSD ($\downarrow$), higher is better for ODG ($\uparrow$).
Model FAD($\downarrow$) LSD($\downarrow$) ODG ($\uparrow$)
375ms 750ms 375ms 750ms 375ms 750ms
AIDD (WavTokenizer 24kHz) 0.042 0.055 0.044 0.085 -2.303 -2.596
AIDD (UniCodec 24kHz) 0.12 0.169 0.049 0.094 -2.753 -3.289
Table 5: FAD, LSD, and ODG (PEAQ) scores across gap sizes for the MAESTRO dataset.
8 Discussion
Latency Analysis.
To further evaluate our method AIDD against CQT Diff+, we compared inference performance and model size. All experiments were conducted on a single NVIDIA A6000 GPU using the MAESTRO test set. Because inference time depends on the number of denoising steps rather than the gap length, we report the average time over both 375 ms and 750 ms gaps.
As shown in Table 6, AIDD (WavTokenizer) achieves the fastest inference time while also being the smallest model and fastest to train.
Model Param size Training time Avg Inference Time (s) Denoising steps
AIDD (WavTokenizer) 90M (81M) 1 day 5.25 1024
AIDD (UniCodec) 90M (210M) 1 day 11.53 1024
CQT Diff+ 242M 4 days 12.54 35
Table 6: Latency comparison of audio-generation, codec type and inverse-model systems.
Limitation.
AIDD has several limitations. First, its performance is constrained by the quality and bandwidth of the underlying tokenizer; among the two single-codebook codecs we tested, WavTokenizer produced smoother transitions and higher fidelity, but reliance on any specific tokenizer still caps the achievable upper bound (see Supplementary Materials).
Second, comparisons with prior work are affected by cross-domain differences: AIDD uses discrete tokens, whereas baselines generate waveforms or spectrograms. Even with recommended settings, mismatches in reconstruction bandwidth and preprocessing may introduce bias, underscoring the need for a unified benchmark spanning token, latent, and continuous models.
Finally, AIDD inherits WavTokenizer’s 24 kHz bandwidth limit, requiring downsampling of higher-rate recordings and restricting outputs to 24 kHz, which may reduce fidelity, unlike other methods that may not subject to that.
Inference Training Gap.
The model may face a training-inference mismatch: during training, the tokenizer sees a fully intact audio signal before masking (tokenize-then-mask), whereas at inference it must tokenize audio that already contains extended gaps (mask-then-tokenize). Although not fixable, We try to quantify this in the Supplementary Material.
@@ 9 Conclusion
We introduced AIDD, a discrete diffusion framework for audio inpainting that operates in a tokenized space using WavTokenizer and a Diffusion Transformer. By modeling discrete tokens rather than raw waveforms, our approach enables stable training and semantically coherent reconstruction, particularly for long gaps. We further proposed span-based masking and a derivative-based regularization loss, which improve structured corruption modeling and temporal smoothness during reconstruction.
Experiments on MusicNet and MAESTRO demonstrate competitive or superior performance, especially for challenging gaps up to 750 ms. Beyond inpainting, our framework establishes a general paradigm for token-based generative modeling in audio. Future work may explore extending these training strategies to other token-based domains, including language models.
@@
Appendix A Information loss Analysis
To assess information loss in WavTokenizer and UniCodec, we evaluate FAD and LSD in two stages. First, we measure loss introduced by the tokenizer by encoding and decoding the original audio and comparing it to the reference. Second, we evaluate inpainting by replacing a 750 ms segment with one treated as if produced by the denoising stage, then stitching it using our standard method.
Results show that reconstruction error in the inpainted case is minimal. This indicates that most quality degradation comes from tokenization, not from the denoising process, highlighting the effectiveness of our method while reflecting its limitations.
Model FAD($\downarrow$) LSD($\downarrow$)
Tokenized Inpainted (750ms) Tokenized Inpainted (750ms)
UniCodec (24kHz) 2.604 0.078 0.498 0.0645
WavTokenizer (24kHz) 1.06 0.061 0.472 0.044
Table 7: FAD and LSD scores across gap sizes for the MAESTRO dataset for tokenization information loss.
@@ Appendix B Reproducibility Statement
With the following hyperparameters (see Table 8) as well as appropriate hardware, anyone can reproducibly conduct our experiment using our code.
@@
Gradient accumulation 1
Training steps (MusicNet / MAESTRO) 100,000 / 150,000
@@
Temp 1.0
top-k 1.0
Span training True
Graph type absorb
Noise schedule loglinear
Predictor Euler
Sampling steps 128
@@ B.1 Training Environment
Appendix C Inference Training Gap Quantization
To further examine this point, we conducted an additional experiment to isolate the effect of masking order at inference. Instead of masking the audio waveform before tokenization (mask-then-tokenize), we first tokenized the clean ground-truth audio and then applied masking directly at the token level, so that the tokenizer operated on uninterrupted audio, matching the training setup. The model then performed inpainting on these masked tokens. The results show that the two procedures achieve very similar performance, with only negligible differences, as shown in Table 9.
Model FAD($\downarrow$) LSD($\downarrow$)
375ms 750ms 375ms 750ms
AIDD (mask-then-tokenize) 0.042 0.055 0.044 0.085
AIDD (tokenize-then-mask) 0.033 0.056 0.040 0.082
Table 9: FAD and LSD scores across gap sizes for the MAESTRO dataset for inference gap analysis.
Appendix D Subjective Evaluation Setup
To assess the perceived audio quality of our method in comparison with competing approaches, we conducted a structured subjective evaluation. The goal of this evaluation was to measure how listeners rate the reconstructed audio when gaps of varying lengths are introduced, and to determine whether our method provides a noticeable improvement in perceptual quality.
We began by selecting the same audio clips used in our objective evaluation, all taken from the MAESTRO dataset test set. For each clip, we inserted a single artificial gap of either 375 ms or 750 ms, matching the conditions used in our quantitative experiments. This ensured consistency between the objective and subjective evaluations and enabled a meaningful comparison across methods.
From the full set of processed audio, we randomly chose five samples for each gap duration (a total of ten samples per method). In addition, we included five unmodified audio segments from the dataset to serve as reference samples representing “perfect quality.” Before presenting the clips to listeners, we anonymized all files by removing any identifying information about the method used for reconstruction. The order of the clips was also randomized to avoid bias and to ensure that participants could not infer patterns.
All selected audio segments were embedded into a Google Forms questionnaire. Participants were instructed to listen to each segment carefully, using headphones if possible, and to rate its overall quality on a 1–5 Likert scale. A score of 5 corresponded to “excellent quality, indistinguishable from the original,” whereas a score of 1 indicated “very poor quality, noticeable artifacts or distortions.”
Once the responses were collected, we aggregated the ratings for each method and each gap length. The results of this analysis, including average scores and comparative rankings, are presented and discussed in the main body of the paper.
We gave the following instructions: You are requested to listen carefully and for each audio rate 1-5 the audio quality in your opinion. 1 being the lowest - very poor quality and 5 being the highest - very high quality. We recommend that you do it on your computer with headsets. Please click on the link, listen, and provide a score based on your opinion.